OpenStudy (anonymous):

SUPER QUICK 4z^-4 is it 1/4z^4 or 4/z^4?

OpenStudy (anonymous):

\[4z ^{-4}\]

OpenStudy (mertsj):

Does the 4 have a negative exponent?

OpenStudy (anonymous):

indeed

OpenStudy (mertsj):

So you really meant that the problem is: \[(4z)^{-4}\]

OpenStudy (anonymous):

not in parenthesis

OpenStudy (mertsj):

So the 4 DOESN'T have a negative exponent.

OpenStudy (anonymous):

\[4z ^{-4}\]

OpenStudy (anonymous):

looks like that

OpenStudy (mertsj):

In that case on what basis could you justify moving it to the denominator?

OpenStudy (mertsj):

\[4z ^{-4}=4^1z ^{-4}=\frac{4}{z^4}\]

OpenStudy (anonymous):

Thanks

OpenStudy (mertsj):

yw

OpenStudy (kainui):

Rather than just blithely giving away the answer to be forgotten later, how do we know that a negative exponent is really just divided like that? Well suppose you have 16/4, that's the same as:\[\frac{ 16 }{ 4} = \frac{ 2^4 }{ 2^2 }\] Now of course we know that 16/4 =4 right? But now we see that we have a really nice rule where:\[\frac{ 16 }{ 4} = \frac{ 2^4 }{ 2^2 }=2^4*2^{-2}=2^{4-2}=2^2=4\] since we already know that exponents add, just play around with exponents and you will see that 2^3*2^2=2^5. we can see that the rule is just subtraction for division. fun stuff